# Range Membership Inference Attacks

### v1

@@ 4.1 Composite hypothesis testing
In this paper, we adopt a simple attack strategy. Our solution to this is based on the type of data in the sampling space. If the data are all naturally ID, we can combine Bayes Factor and GLRT by taking the average likelihood of the top samples to reduce the influence of the randomness in the sampling process. On the other hand, if the adversary can only synthesize data within the range, the top samples are highly likely to be OOD data with high membership scores. Hence, in this case, we want to remove those points from the equation. Since the presence of training points intuitively raises the average membership score of ID points nearby, compared to having no training points at all, we average the membership scores of the remaining samples. Unifying these two strategies gives us the following:
@@
$\mathbb{P}\left(\theta|H_{1}\right)=\text{TrimmedAvg}(S,q_{s},q_{e};\mathbb{P})=\text{Avg}_{x\not\in[q_{s},q_{e}]\text{-th quantiles}}\mathbb{P}(\theta|x\in D),$ (1)
@@
where $S$ is the sampled set, $q_{s}$ and $q_{e}$ mark the start and the end of the quantiles where we want to remove to compute our robust statistics that are one-sided trimmed means. If the sampling space is filled with synthetic data, the chance of the top samples being false positives is high, so we set $q_{e}=100$ to remove the largest points. $q_{s}$ is a hyperparameter that decreases (trim more) as the quality of sampled points gets worse. On the other hand, if the sampling space consists of real points, we set $q_{s}=0$ to remove the smallest points in our aggregation. $q_{e}$ decreases (trim less) as the number of real samples decreases to offset the high variance due to limited samples available. Note that the optimal hyperparameters may differ across different membership signals (e.g. loss values, LiRA scores), as they exploit different vulnerabilities and expose different training points. However, for fixed model architectures, range functions, data distributions and sampling methods, these hyperparameters can be determined by reference models, similar to the offline version of RMIA [47]. Specifically, by randomly choosing a reference model as the temporary target model, we can run RaMIAs using all other reference models while sweeping these hyperparameters.
@@ 4.2 Range membership inference attack as a framework
The range membership inference attack is a new inference attack framework, not a particular attack algorithm. There are two components in this framework: a sampler and a membership tester, both of which are necessary to compute the range membership score formulated in Eqn 1. The sampler $\text{Sampler}(\mathcal{R}):S\rightarrow X$ returns samples within the given range. The membership tester $\text{MIA}(x)$ is a (point-query) membership inference algorithm that outputs a membership score, which can be used to approximate $\mathbb{P}(\theta|x)$. A number of existing attack algorithms can be plugged in. Similar to MIAs, the key to using RaMIA as a privacy auditing tool is to compute the range membership score. Our framework can adopt any existing membership scoring function $\text{MIA}(x)$ to compute $\text{RaMIA}(\mathcal{R})$. Below, we outline the attack with our attack strategy described above:
@@ 5 Experiments
Since the purpose of this paper is to introduce a new concept and framework, the goal of the experiments section is to provide a proof-of-concept. We experiment on the commonly used Purchase-100 [39], CelebA [28], CIFAR-10 [25] and AG News [48] datasets. Details of how we split the dataset, train models, construct ranges, and obtain samples are explained in Appendix B. We compare RaMIA with MIA (Sec 5.1) in scenarios MIA under-performs (Depicted in Figure 2). Both attacks are built upon the state-of-the-art attack algorithm, robust membership inference attack (RMIA) [47], with three reference models trained in the same way as Carlini et al. [5] and Zarifzadeh et al. [47]. The respective queries are outlined in Table 1, and the definitions of members under each attack framework are explained in Table 2. In both tables, $x$ represents original data in datasets, while $x^{\prime}$s are either data with missing values or modified data from $x$. The reason that we do not test our attacks by taking ranges centered at original data $x$ is that the chance of the attack data being exactly the same as the training data is extremely low without sufficient prior knowledge. It is more realistic that similar data are being queried.
@@
Table 2: Definitions of range and point members corresponding to the attack queries in Table 1
@@
On Purchase-100, we take 20 samples in every range, and set $q_{e}=100,q_{s}=45$. On CIFAR-10, we apply up to 15 distinct transforms, and set $q_{e}=100,q_{s}=40$. On AG News, we construct 50 sentences within each range, and set $q_{e}=100,q_{s}=20$. On CelebA, each celebrity has a different number of images in the sampling space, ranging from 1 to 18. Since it is hard to standardize the sample size for all ranges, we take all of them. We then set $q_{s}=0$ and $q_{e}=25$, which means we are not trimming anything for ranges with very few samples available.
5.1 RaMIAs quantify privacy risks more comprehensively than MIAs
As we have explained before, data points that are close enough to the training data are out of the scope of membership inference attacks. We observe from Figure 3 that range membership inference attacks are better at identifying those nearby points, and thus providing more comprehensive privacy auditing on all the four datasets we tested. We want to emphasize that the gain is remarkable if we consider how little samples were taken compared to the range sizes. On Purchase-100, there are a total of 1024 candidates, and we take less than 20% of them. On AG News, there are millions of sentences within a distance of 8. 50 sentences are too little to meaningfully cover anything in the space. But yet limited samples can lead to noticeable gains, which further shows the current privacy quantification approach is flawed and needs a better framework. Due to randomness in sampling, we report the average gain of RaMIA over MIA with standard deviation in Table 4.TPRs at small FPRs are in Table 3
@@ 5.1 RaMIAs quantify privacy risks more comprehensively than MIAs
5.2 Factors affecting RaMIA performance
Training data density in the range
(b) RaMIA performs better on CIFAR-10 when the range size increases. The size is equal to the number of distinct transformation applied to images.
Susceptibility to MIAs and RaMIAs is correlated
Ranges containing training points that are susceptible to MIAs are also more susceptible to RaMIAs. Researchers have previously discovered that machine learning models memorize duplicate data more [26, 6]. In our CelebA dataset, each celebrity has different numbers of photos in the training set, which can be thought as that each identity has different levels of duplication in the training set. Similar to the insights from MIAs, we also observe that identities that have more training images, i.e. higher duplication rate, are more susceptible to RaMIA. Figure 8 shows the relationship between the percentile each range’s RaMIA score within non-members’ RaMIA scores and the duplication rate. Generally speaking, identities that have more training photos are more prone to RaMIAs. Similarly correlation can be observed on the other three datasets in our experiments, where the training points’ RaMIA score percentiles among non-members are positively correlated with their MIA score percentiles 7.
5.3 Mismatched training and attack data hurts attack performance
Difference to existing augmentation-based MIAs
Existing attacks [4, 47] also use augmented queries in the attack, but with a different rationale and assumption of the attacker’s knowledge. Since they adopt the existing privacy notion based on point queries, only (augmented) images seen by the model in the training stage are considered as members. Hence, the attacker needs to know the exact train-time augmentations and augment images accordingly to not violate the privacy of notion. In RaMIA, the set of augmentations is given by the challenger (Def 2), which can contain augmentations not used in training, but considered as privacy leaking. Using the aggregation method in [4] will hurt the attack performance if non-training augmentations are used. However, RaMIA is designed to be robust in this scenario (Fig 3(c)).
6 Conclusion
In this paper, we argue that membership inference attacks are only useful as a privacy audit tool when querying exact copies of training and test data. Moving the query to similar points causes a drastic decrease in performance, rendering MIAs less useful. We conclude MIAs fail to comprehensively capture the notion of privacy, and thus propose a new class of inference attack, RaMIA, that extends the notion of MIAs. and cover the failure cases of MIAs by checking if a given range contains a training point. We introduce RaMIA as an attack framework that can be implemented with any existing MIA algorithm. We show that it can provide better privacy auditing with very few samples taken randomly. We hope our work can make more privacy researchers and practitioners aware of the shortcomings of MIAs, and shift their attention to RaMIAs. As it is the first paper that brings up this new framework, there is room for improvement in specific attack algorithms. For example, a better sampling process will surely increase the gap between RaMIA and MIA. Nevertheless, we have shown our framework is sensible and useful. In future work, we hope to design more powerful RaMIA strategies that are robust to the change of membership signals and datasets, especially on LLMs where we believe our privacy notion is extremely relevant.
Appendix A Attack Details
A.1 (Simple) Hypothesis testing
$H_{0}$ $:\text{The given }z\text{ is not a training point }(b=0).$
$H_{1}$ $:\text{The given }z\text{ is a training point }(b=1).$
The likelihood ratio test (LRT) is then conducted
$\frac{\mathbb{P}\left(\theta|H_{1}\right)}{\mathbb{P}\left(\theta|H_{0}\right)}$ (2)
This is usually called "simple" hypothesis testing because each $H$ contains a single hypothesis.
A.2 Attack algorithms
LOSS
LOSS [45] computes loss values as a proxy of membership score on given points: $\text{MIA}(x;\theta)=\mathit{l}(x;\theta)$. To compute the likelihood, an easy way is to take the exponential of the negative of the loss $P=\exp^{-\mathit{l}}$.
RMIA
RMIA [47] computes membership score by applying chain rule: $\mathbb{P}(\theta|x)=\frac{\mathbb{P}(x|\theta)\mathbb{P}(\theta)}{\mathbb{P}(x)}$. The score is then compared with all available population data points to obtain the percentage of population points being dominated by the given point: $\mathbb{P}_{z\in Z}[\frac{\mathbb{P}(\theta|x)}{\mathbb{P}(\theta|z)}\leq\gamma]$, where the term $\mathbb{P}(\theta)$ will cancel out with each other. The normalizing constant $\mathbb{P}(x)$ is computed with reference models: $\mathbb{P}(x)=0.5\mathbb{E}_{\theta_{\text{in}}}\mathbb{P}(x|\theta_{\text{in}})+0.5\mathbb{E}_{\theta_{\text{out}}}\mathbb{P}(x|\theta_{\text{out}})$. In its offline version, the in models are unavailable. In this case, the former probabilities is approximated by the latter term $\mathbb{P}_{\text{in}}=a\mathbb{P}_{\text{out}}+(1-a)$. The hyperparameter $\alpha$ is chosen based on the reference models. Specifically, one reference model is chosen as the temporary target model, and the rest are used to attack it. The value of $\alpha$ is chosen to be the best performing value under this setting, obtained via a simple sweeping. In our experiment, we use the offline attack only. The $\alpha$ values for Purchase-100 and CIFAR-10 are taken from [47]. For CelebA, we set it to be 0.33. For AG News, we set it to be 1.0.
Appendix B Setup Details
On each of the dataset, we train four models on half of the dataset in the same way described by Carlini et al. [5], Zarifzadeh et al. [47]. We will describe the details of the datasets below. We have checked with their licenses with our best effort, and confirm their terms of use are respected.
B.1 Tabular data: Purchase-100
Dataset
Purchase-100 [39] is a tabular dataset derived from Kaggle’s Acquire Valued Shoppers Challenge  [^1] https://www.kaggle.com/c/acquire-valued-shoppers-challenge/data. This dataset was first curated by Shokri et al. [39] such that there are 600 binary features, each representing if each person, represented by each row, has purchased the product. The data is then divided into 100 classes, and the task is to predict the category of the person given the purchase history.
Models
We train five-layer multi-layer perceptron (MLP) models in PyTorch [33] on half of the entire dataset. The hidden layers are of sizes $[512,256,128,64]$. All models achieve a test accuracy of $83\%$.
Construction of ranges
We simulate the scenario where the attacker has incomplete data (data with missing values). For all training and test data records, we randomly mask k columns. Each row with masked columns is a range query that contains $2^{k}$ possible points as each feature is binary. We then check if each range constructed from test data points include any training point. If so, they are re-labelled as "in-ranges".
Sampling within ranges
Since this dataset contains 600 independent binary features, we do Bernoulli sampling independently for all missing columns. The parameter of the sampler is computed by taking the average value of each column. Because of the nature of this dataset, our sampled data can be regarded as in-distribution. We take 19 samples for each range, together with the data obtained by doing mode imputation (fill in the missing values with the modes).
B.2 Image data I: CelebA
Dataset
CelebA [28], also known as the CelebFaces Attributes dataset, contains 202,599 face images from 10,177 celebrities, each annotated with 40 binary facial features. We construct the members set by only including photos of celebrities with identity number smaller than 5090. The rest are used to construct the non-members set. For each celebrity in the members set, half of the photos are put into the training set, while the other half goes into the holdout set.
Model
We train four-layer convolutional neural networks (CNNs) in PyTorch [33] on the training set to predict the facial attributes of any given photo. Our target model has a test accuracy of $87\%$.
Construction of ranges
The range function here is a semantic one that is based on the identity of the face image. For example, a range query can be "all Alice’s photos". Since the identities in the training and non-members set are disjoint, it is easy to construct in- and out-ranges based on the distribution of identities in the two sets.
Sampling within ranges
For each range query, we curate all images in the holdout set that share the same identity as the range center to construct our sample set.
B.3 Image data II: CIFAR-10
Dataset
CIFAR-10 [25] is a popular image classification dataset. There are 50,000 training images, each of size $(32,32,3)$.
Models
We train WideResNets-28-2 [46] with JAX [2] on half of the training set of CIFAR-10 using the code from [5], with and without image augmentations. Our target model trained without augmentation achieves a test accuracy of $83\%$ on the CIFAR-10 test set, and the target model trained with augmentation achieves a test accuracy of $92\%$. The train time augmentation is the composition of random flipping, cropping and random hue.
Construction of ranges
The range function here is different types of image augmentations, which are geometric transformations. An example of a range query is "all transformed version of image $X$". For each training and test image, a range is constructed by applying different transformations.
Sampling within ranges
For each range query, we independently apply 15 image augmentations on the query center. The augmentations include flipping, random rotation, random resizing and cropping, random contrast, brightness, hue, and the composition of them.
B.4 Textual data: AG News
Dataset
@@ B.4 Textual data: AG News
We use AG News dataset [48], which is a news collection with four categories of news. Although it was introduced as a text classification dataset, we disregard the labels and treat it as a text generation dataset. There are 120,000 sentences in its training set.
Model
@@
We took pretrained GPT-2 [34] models from Hugging Face’s transformers library, and finetuned them on half of AG News’ training set with LoRA [19] implemented in Hugging Face’s PEFT [32] library. The finetuning is done for 4 epochs. Our target model achieves a perplexity of 1.39 on the test set of AG News.
Construction of ranges
@@
The range function here is word-level Hamming distance, which can be thought as the edit distance measured on word level that only allows word substitution. An example of a range query is "all sentences within Hamming distance $d$ to sentence $x$". To construct in- and out-ranges, we just need to specify the max Hamming distance and the starting sentence. We constructed the starting sentences by randomly masking $\alpha$ words from the training and test sentences, before filling in the mask with a pre-trained BERT [12] model, so they have a distance of $\alpha$ to the original training/test sentences. A Hamming distance is then specified with each starting sentence to form a range.
Sampling within ranges
@@
We mask the range center by $k$ words where $k$ is the Hamming distance specified by the range. Then we use BERT [12] to replace the mask with one of the top choices.
Appendix C Implementation details
@@
For all PyTorch models, we use Adam as our optimizer with a learning rate of 0.001. For WideResnets, we use the training code from [4]. On AG News, the models are trained for 4 epochs. On other datasets, they are trained for 100 epochs.
All training are done on two Nvidia RTX 3090 GPUs. Training on AG News takes about 1 hour per epoch. Training other models takes less than one hour each.
Appendix D RaMIA on redacted data
Many large language models (LLMs) are trained with sensitive textual data. Some of the data with sensitive information redacted might be public available. Similar to our experiment with data with missing values, we can apply RaMIA to redacted data to identify which of them are used to train a target LLM. Accurately identify the redacted sentences paves the way for reconstructing them as a follow-up attack. Figure 5 shows the results. In this experiment, we use spaCy [18] to mask peoples’ names to simulate masking of personally identifiable information (PII). We then generate 10 possible sentences for each masked sentence using BERT, and conduct RaMIA. The MIA performance is the average attack performance over all 10 possible sentences.
Figure 5: RaMIA on subset of AG News where names are redacted.
Appendix E Extra results
In this section, we put extra experiment results.

### v2

@@ 4.3 Our approach: Trimmed averages
Given these limitations, in this paper, we propose a simple attack strategy based on Bayes Factor. We first assume every sampled point is of equal probability. Then, to address the unreliability of MIA scores, we use a trimming process based on the type of data in the sampling space. If the data are all naturally ID, we trim the less probable samples and take the average likelihood of the top samples to reduce the influence of non-members and the randomness in the sampling process. On the other hand, if the adversary can only synthesize data within the range instead of sampling from a pool of real data, the top samples are highly likely to be OOD data with high membership scores. Hence, in this case, we want to remove those points from the equation. We then trim the top-scoring samples and take the average of the rest. This gives us the following:
@@
$\end{multlined}\text{I\kern-1.49994ptP}{\left(\theta|H_{1}\right)}=\text{TrimmedAvg}(S,q_{s},q_{e};\text{I\kern-1.49994ptP})\\ =\text{Avg}_{x\not\in[q_{s},q_{e}]\text{-th quantiles}}\text{I\kern-1.49994ptP}{(\theta|x\in D)},$ (4)
@@
where $S$ is the sampled set, $q_{s}$ and $q_{e}$ mark the start and the end of the quantiles that we want to remove to compute our robust statistics that are one-sided trimmed means. If the sampling space is filled with synthetic data, the chance of the top samples being false positives is high, so we set $q_{e}=100$ to remove the largest points. $q_{s}$ is a hyperparameter that decreases (trim more) as the quality of sampled points gets worse. On the other hand, if the sampling space consists of real points, we set $q_{s}=0$ to remove the smallest points in our aggregation. $q_{e}$ decreases (trim less) as the number of real samples decreases to offset the high variance due to limited samples available. If we do not trim anything, the formulation reduces to Bayes Factor which assigns equal probability to all points.
Note that the optimal hyperparameters may differ across different membership signals (e.g. loss values, LiRA scores), as they exploit different vulnerabilities and expose different training points. However, for fixed model architectures, range functions, data distributions, and sampling methods, these hyperparameters can be determined by reference models, similar to the offline version of RMIA [48]. Specifically, by randomly choosing a reference model as the temporary target model, we can run RaMIAs using all other reference models while sweeping these hyperparameters in a grid search manner. More details are in Appendix A.
@@ 4.4 Range membership inference attack as a framework
The range membership inference attack is a new inference attack framework, not a particular attack algorithm. There are two components in this framework: a sampler and a membership tester, both of which are necessary to compute the range membership score formulated in Eqn 4. The sampler $\text{Sampler}(\mathcal{R}):S\rightarrow X$ returns samples within the given range. The membership tester $\text{MIA}(x)$ is a (point-query) membership inference algorithm that outputs a membership score, which can be used to approximate $\text{I\kern-1.49994ptP}{(\theta|x)}$. Any existing MIA algorithm can be plugged in as RaMIA’s membership testing backbone. Similar to MIAs, the key to using RaMIA as a privacy auditing tool is to compute the range membership score. Our framework can adopt any existing membership scoring function $\text{MIA}(x)$ to compute $\text{RaMIA}(\mathcal{R})$. Below, we outline the attack with our attack strategy described above:
@@ 5 Experiments
Since the purpose of this paper is to introduce a new concept and framework, the goal of the experiments section is to provide a proof-of-concept. We experiment on the commonly used Purchase-100 [40], CelebA [28], CIFAR-10 [25] and AG News [49] datasets. We will describe the details of how we split the dataset, train models, construct ranges, and obtain samples in Section 5.1. Since the range membership notion is new, we do not have a prior method to compare with. However, since the goal of introducing this new privacy notion is to enable better and more comprehensive privacy auditing, we compare RaMIA with the de-factor privacy auditing framework, MIA. We show that RaMIA performs better (Sec 5.5) in scenarios where MIA under-performs (Depicted in Figure 2). Both MIA and RaMIA in the experiments are built upon the state-of-the-art attack algorithm, robust membership inference attack (RMIA) [48], with a few reference models trained in the same way as [5] and [48]. The respective queries are outlined in Table 1, and the definitions of members under each attack framework are explained in Table 2. In both tables, $x$ represents original data in datasets, while $x^{\prime}$s are either data with missing values or modified data from $x$. The reason that we do not test our attacks by taking ranges centered at original data $x$ is that the chance of the attack data being exactly the same as the training data is extremely low without sufficient prior knowledge. It is more realistic that similar data are being queried.
@@
Table 2: Definitions of range and point members corresponding to the attack queries in Table 1.
@@
5.1 Setup details
As mentioned earlier, the range function must come from the semantics of privacy. Hence, in experiments, we use specific range functions with different data types. For tabular data, we consider missing columns, which is an extreme case of using Euclidean distance on missing columns as a range function. For human photo data, we use a semantic range function, which is based on the main person featured in the photo. For other image data, we use geometric transformations as range functions. For text data, we use (word level) Hamming distance, which is edit distance or Levenshtein distance that only considers word substitution. The reasons for choosing these range functions have been motivated and explained in earlier parts of this paper.
5.1.1 Tabular data: Purchase-100
  • Dataset Purchase-100 [40] is a tabular dataset derived from Kaggle’s Acquire Valued Shoppers Challenge  [^1] https://www.kaggle.com/c/acquire-valued-shoppers-challenge/data. This dataset was first curated by [40] such that there are 600 binary features, each representing if each person, represented by each row, has purchased the product. The data is then divided into 100 classes, and the task is to predict the category of the person given the purchase history.
  • Models We train five-layer multi-layer perceptron (MLP) models in PyTorch [34] on half of the entire dataset. The hidden layers are of sizes $[512,256,128,64]$, identical to the setup in [48]. All models achieve a test accuracy of $83\%$.
  • Construction of ranges We simulate the scenario where the attacker has incomplete data (data with missing values). For all training and test data records, we randomly mask k columns. Each row with masked columns is a range query that contains $2^{k}$ possible points as each feature is binary. We then check if each range constructed from test data points includes any training point. If so, they are re-labeled as ”IN-ranges”.
  • Sampling within ranges Since this dataset contains 600 independent binary features, we do Bernoulli sampling independently for all missing columns. The parameter of the sampler is computed by taking the average value of each column. Because of the nature of this dataset, our sampled data can be regarded as in-distribution. We take 19 samples for each range, together with the data obtained by doing mode imputation (fill in the missing values with the modes).
5.1.2 Image data I: CelebA
  • Dataset CelebA [28], also known as the CelebFaces Attributes dataset, contains 202,599 face images from 10,177 celebrities, each annotated with 40 binary facial features. We construct the members set by only including photos of celebrities with identity numbers smaller than 5090. The rest are used to construct the non-members set. For each celebrity in the members set, half of the photos are put into the training set, while the other half goes into the holdout set.
  • Models We train four-layer convolutional neural networks (CNNs) in PyTorch [34] on the training set to predict the facial attributes of any given photo. Our target model has a test accuracy of $87\%$.
  • Construction of ranges The range function here is a semantic one that is based on the identity of the face image. For example, a range query can be ”all Alice’s photos”. Since the identities in the training and non-members set are disjoint, it is easy to construct IN- and OUT-ranges based on the distribution of identities in the two sets.
  • Sampling within ranges For each range query, we curate all images in the holdout set that share the same identity as the range center to construct our sample set.
5.1.3 Image data II: CIFAR-10
  • Dataset CIFAR-10 [25] is a popular image classification dataset. There are 50,000 training images, each of size $(32,32,3)$.
  • Models We train WideResNets-28-2 [47] with JAX [2] on half of the training set of CIFAR-10 using the code from [5], with and without image augmentations. Our target model trained without augmentation achieves a test accuracy of $83\%$ on the CIFAR-10 test set, and the target model trained with augmentation achieves a test accuracy of $92\%$. The train time augmentation is the composition of random flipping, cropping and random hue.
  • Construction of ranges The range function here is different types of image augmentations, which are geometric transformations. An example of a range query is ”all transformed version of image $X$”. For each training and test image, a range is constructed by applying different transformations.
  • Sampling within ranges For each range query, we independently apply 15 image augmentations on the query center. The augmentations include flipping, random rotation, random resizing and cropping, random contrast, brightness, hue, and the composition of them.
@@ 5.1.3 Image data II: CIFAR-10
5.1.4 Textual data: AG News
@@
  • Dataset We use the AG News dataset [49], which is a news collection with four categories of news. Although it was introduced as a text classification dataset, we disregard the labels and treat it as a text generation dataset. There are 120,000 sentences in its training set. It is a popular text dataset in LLM MIA literature [33].
@@
  • Models We took pretrained GPT-2 [35] models from Hugging Face’s transformers library, and finetuned them on half of AG News’ training set with LoRA [19] implemented in Hugging Face’s PEFT [32] library. The finetuning is done for 4 epochs. Our target model achieves a perplexity of 1.39 on the test set of AG News.
@@
  • Construction of ranges The range function here is word-level Hamming distance, which can be thought of as the edit distance measured on word level that only allows word substitution. An example of a range query is ”all sentences within Hamming distance $d$ to sentence $x$”. To construct IN- and OUT-ranges, we just need to specify the max Hamming distance and the starting sentence. We constructed the starting sentences by randomly masking $\alpha$ words from the training and test sentences, before filling in the mask with a pretrained BERT [12] model, so they have a distance of $\alpha$ to the original training/test sentences. A Hamming distance is then specified with each starting sentence to form a range.
@@
  • Sampling within ranges We mask the range center by $k$ words where $k$ is the Hamming distance specified by the range. Then we use BERT [12] to replace the mask with one of the top choices.
5.2 Hyperparameters
Overall, on Purchase-100, we take 20 samples in every range, and set $q_{e}=100,q_{s}=45$. On CIFAR-10, we apply up to 15 distinct transforms, and set $q_{e}=100,q_{s}=40$. On AG News, we construct 50 sentences within each range, and set $q_{e}=100,q_{s}=20$. On CelebA, each celebrity has a different number of images in the sampling space, ranging from 1 to 18. Since it is hard to standardize the sample size for all ranges, we take all of them. We then set $q_{s}=0$ and $q_{e}=25$, which means we are not trimming anything for ranges with very few samples available.
5.3 Implementation details
@@
We train 8 models on Purchase-100 and CIFAR-10, and train 4 models on CelebA and AG News. Each model is trained on half of the dataset, as described in [5, 48]. For all PyTorch models, we use Adam as our optimizer with a learning rate of 0.001. For WideResnets, we use the training code from [5]. On AG News, the models are trained for 4 epochs. On other datasets, they are trained for 100 epochs. All training is done on two Nvidia RTX 3090 GPUs. Training on AG News takes about 1 hour per epoch. Training other models takes less than one hour each.
5.4 Metrics
To evaluate the performance of RaMIA and MIA based on the same MIA algorithm, we use AUCs. The inputs to RaMIA are range queries in Table 1, while the inputs to MIA are point queries, which are range centers, in Table 1. The ground truth membership for both RaMIA and MIA is the range membership information defined in Table 2, which allows us to compare the power of both inference attack frameworks in identifying queries that are privacy leaking. In other words, both point and range TPRs/FPRs are calculated against the range membership information. Following [5], we also report (Range) TPR at small (Range) FPR for both MIA and RaMIA (Table 4).
5.5 RaMIAs quantify privacy risks more comprehensively than MIAs
As we have explained before, data points that are close enough to the training data are out of the scope of membership inference attacks. We observe from Figure 3 that range membership inference attacks are better at identifying those nearby points, and thus providing more comprehensive privacy auditing on all the four datasets we tested. We want to emphasize that the gain is remarkable if we consider how little samples were taken compared to the range sizes. On Purchase-100, there are a total of 1024 candidates, and we take less than 20% of them. On AG News, there are millions of sentences within a distance of 8. 50 sentences are too little to meaningfully cover anything in the space. Yet, limited samples can lead to noticeable gains, which further shows the current privacy quantification approach is suboptimal and needs a better framework. Due to randomness in sampling, we report the average gain of RaMIA over MIA with standard deviation in Table 3.TPRs at small FPRs are in Table 4. The improvement in AUC is summarized in Table 3.
Table 3: Improvement in AUCs after switching from MIA to RaMIA across multiple iterations of random sampling. We do not randomly sample but use all available attack images in CelebA.
Purchase-100 CIFAR-10 CelebA AG News
$\Delta\text{AUC}$ $2.62\pm 0.04$ $4.12\pm 0.06$ $5.4$ $1.20\pm 0.2$
Relation to user-level inference
Note that in the CelebA experiment, we use identify information as the semantic range function, making it similar to user-level inference. This further shows that RaMIA is a better and more comprehensive privacy auditing framework. In terms of algorithms, our attack strategy should dominate the simple averaging approach used in prior work [31, 23], since the trimming ratio is optimized (compared to no trimming in simple averaging). There are other user-level inference algorithms [11] that train additional shadow models and classifiers, which incurs additional computational costs. They also consider scenarios where user information is the label of their target models, e.g. facial and speech recognition systems that predict user ID as their outputs [11, 10]. This encourages their target models to explicitly cluster data based on user information, making inference easier. On the other hand, our target model, a facial attribute classifier, does not use user information anywhere in the training, making the inference harder.